A Smoothie king on a large college campus recently sold small low-fat fruit smoothies for $4.30 each, medium smoothies for $6.50 each, and large smoothies for $8.00 each. One hot summer afternoon, Jake sold 34 smoothies for a total of $211. The number of small and large smoothies, combined, was 2 more than the number of medium smoothies. How many of each size were sold?
X= # small smoothies
Y=# medium smoothies
Z=# large smoothies
X+Y+Z=34
X+Z=2X+Y
4.30X+6.50Y+8.00Z=211

Answer by richwmiller(17219)

X+Y+Z=34
X+Z=Y+2 [not X+Z=2X+Y]
4.30X+6.50Y+8Z=211
X=10., Y=16., Z=8.